OpenStudy (chacha1234):

Please help with this quick 8th grade math question? Will medal and fan! At 3:30 p.m., Berto’s train was 34 miles past the egg farm, traveling at an average speed of 85 miles per hour. At the same time on a nearby track, Eduardo’s train was traveling at an average speed of 110 miles per hour and had 12 miles to go before it reached the egg farm. To the nearest hundredth of an hour, after how much time will the trains meet up?

OpenStudy (chacha1234):

Here are the choices: 0.11 hours 0.25 hours 0.88 hours 1.84 hours

OpenStudy (chacha1234):

@RhondaSommer

OpenStudy (moldybubblegum12):

.25 hours

OpenStudy (chacha1234):

Thanks

OpenStudy (chacha1234):

@moldybubblegum12 One more?

OpenStudy (moldybubblegum12):

ok

OpenStudy (chacha1234):

Which system of equations below has no solution? y = 4x + 5 and y = 4x – 5 y = 4x + 5 and 2y = 8x + 10 y = 4x + 5 and y = 1/4x + 5 y = 4x + 5 and y = 8x + 10

OpenStudy (moldybubblegum12):

i don't know the answer to this question, maybe this site will help? https://www.wyzant.com/resources/answers/115245/how_to_solve_the_system_of_equations_for_4x_y_5_and_8x_5_2y

OpenStudy (chacha1234):

What about this one @moldybubblegum12

OpenStudy (chacha1234):

What is the solution to the system of equations below? y=-1/2x-21 y=-6x-10

OpenStudy (chacha1234):

You there?

OpenStudy (moldybubblegum12):

yes, honestly i really dont know this one D:

OpenStudy (moldybubblegum12):

Maybe this site will help? http://www.geteasysolution.com/y=-2x-21

RhondaSommer (rhondasommer):

\[\frac{-1}{2}x-21=-6x-10\]\[\frac{-1}{2}x=-6x+11\]\[5\frac{1}{2}x=11\]\[x=2\]\[-6(2)-10\]\[-12-10\]\[y=-22\]

RhondaSommer (rhondasommer):

(2,-22) is your answer